OpenStudy (arbyj):

Someone me help solve this; 4a+3=60a-45 I have the answer but I want to make sure its correct (Please show work)

Directrix (directrix):

@arbyj Do you have the work that gave the answer? If so, please post it for checking. Thanks.

OpenStudy (arbyj):

I subtracted 5 from both sides and got this -6x+12≥33 then -12 -12 -6x ≥ 21 6 6 x ≤ -3.5 @Directrix

OpenStudy (anonymous):

start by subtracting 5 from both sides\[-3(2x-4)\ge33\] then divide by (-3) switching the inequality\[(2x-4)\le-11\] add 4 \[(2x)\le-7\] divide by 2\[x \le-3.5\]

Directrix (directrix):

Yes, that is correct. @arby